Of the members of a Boy scout troop, 15% have received the First Aid merit badge. If 40 Boy Scouts are selected at random, use the normal distribution as an approximation to the binomial distribution to find the probability that four or more will have the First Aid merit badge.
MULTIPLE CHOICE:
A)86.6%
B)81.3%
C)36.6%
D)31.3%

Answer by stanbon(75887)

Of the members of a Boy scout troop, 15% have received the First Aid merit badge. If 40 Boy Scouts are selected at random, use the normal distribution as an approximation to the binomial distribution to find the probability that four or more will have the First Aid merit badge.
------
u = np = 0.15*40 = 6
s = sqrt(6*0.85) = 2.258
-------------------------------
z(3.5) = (3.5-6)/[2.258] = -1.1072
-----
P(4<= x <=15) = P(-1.1072<= z <= 100) = 0.8659
-------------
Cheers,
Stan H.
===============
MULTIPLE CHOICE:
A)86.6%
B)81.3%
C)36.6%
D)31.3%